# InjecGuard: Benchmarking and Mitigating Over-defense in Prompt Injection Guardrail Models

### v2

@@ 5.1 Experimental Setups
Evaluation Datasets. As mentioned in Sec. 3.3, we conduct our evaluations based on three aspects: benign accuracy, malicious accuracy, and over-defense accuracy. To fully verify the practicality of our proposed InjecGuard, we ensure that all evaluation datasets are not included in the training data. Specifically, For evaluating the benign accuracy, we utilize benign data from both the PINT benchmark LakeraAI (2024b) and WildGuard benchmark Han et al. (2024). For evaluating the malicious accuracy, We use attack data from both the PINT benchmark LakeraAI (2024b) and BIPIA datasets Yi et al. (2023). For evaluating the over-defense accuracy, we use the proposed NotInject dataset.
Metrics. Since prompt guard models function as text classification systems that predict whether an input text is benign or malicious, we evaluate their performance using Accuracy, calculated as the proportion of correct predictions over the total number of test cases in the evaluation dataset:
Accuracy=Number of Correct PredictionsTotal Number of Test $\text{Accuracy}=\frac{\text{Number of Correct Predictions}}{\text{Total Numberof Test Cases}}$ (1)
Additionally, we report the computational overhead in terms of Giga Floating Point Operations (GFLOPs), which quantifies the total number of floating-point operations required during inference. GFLOPs provide an estimate of the computational resources needed by the model. We also measure the inference time to assess the computational overhead introduced by the models.
@@ 5.2 Main Results
We compare our method with other baselines in terms of performance and overhead. The results are shown in Tab. 2 (detailed results for each subsets deferred to Appendix. B.1).
Performance Comparison. Our InjecGuard demonstrates superior performance compared to existing prompt guard models and even rivals commercial LLMs like GPT-4o. Specifically, InjecGuard achieves an average accuracy of 83.48%, the best-performing commercial prompt guard model, by 6.25%, and exceeding the top open-source model, ProtectAIv2, by 30.83%. Our model excels across all evaluation categories, attaining an over-defense accuracy of 87.32%, benign accuracy of 85.74%, and malicious accuracy of 77.39%. This balanced performance indicates that InjecGuard is highly effective at correctly identifying benign and malicious inputs while minimizing false positives and negatives. The fact that InjecGuard achieves results comparable to GPT-4o, despite being based on the open-source data and lightweight backbone, highlights the efficiency and effectiveness of our training approach. Notably, according to Tab. 2, the Efficiency (Performance/Inference time) of our model is 503 times higher than that of GPT-4o.
@@
Overhead Comparison. Our model, InjecGuard, excels in both computational efficiency and performance compared to existing baseline models. As shown in Table 2, InjecGuard  achieves an average accuracy of 83.48% while maintaining a GFLOPS count of 60.45 and an inference time of only 15.34 milliseconds. With an efficiency score of 5.44 (performance divided by inference time), InjecGuard delivers robust results swiftly and resource-effectively, making it ideal for environments prioritizing both accuracy and efficiency. In contrast, LLMs like GPT-4o achieve slightly higher accuracy (85.53%) but at a significant computational cost, with an inference time of 7907.18 milliseconds and an efficiency score of 0.01. Similarly, models such as Llama-2-chat and LlamaGuard3 require thousands of GFLOPS and longer inference times, resulting in much lower efficiency. This disparity highlights the impracticality of deploying these LLMs in scenarios where computational resources and quick response times are essential. Therefore, InjecGuard presents a highly effective and efficient alternative for real-world applications demanding both high accuracy and operational efficiency.
Training Design Over-defense (%) Benign (%) Malicious (%) Average (%)
Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
   + MOF with Retraining from Scratch (Sec. 4.2) 89.38 84.73 71.57 81.89
Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
   + MOF with Finetuning (Sec. 4.2) 68.14 82.11 74.81 75.02
Training on the Basic Dataset (Sec. 4.1) 75.22 78.53 70.17 74.64
   + Data-centric Augmentation (Sec. 4.1) 64.31 81.36 75.95 73.87
   + MOF with Retraining from Scratch (Sec. 4.2) 87.32 85.74 77.39 83.48
Table 3: Ablation study of each training design introduced in Sec. 4.
@@ 5.3 Ablation Studies
The ablation study presented in Tab. 3 offers key insights into the effects of our training components on InjecGuard performance. Starting with the Basic Dataset, the model establishes a baseline average accuracy of 74.64%, with over-defense, benign, and malicious accuracies of 75.22%, 78.53%, and 70.17%, respectively. Introducing Data-centric Augmentation alone leads to a noticeable improvement in benign accuracy (from 78.53% to 81.36%) and malicious accuracy (from 70.17% to 75.95%), enhancing the model’s ability to correctly classify both benign and malicious inputs. However, this augmentation comes at a cost, as evidenced by a significant reduction in over-defense accuracy from 75.22% to 64.31%. This decline indicates that while data-centric augmentation enriches the training data and improves classification capabilities, it inadvertently exacerbates the over-defense issue, making the model more prone to incorrectly classifying benign inputs as malicious—a challenge commonly observed in existing models.
@@ 6 Conclusions
In this study, we delved into the over-defense phenomenon observed in existing injection-based prompt guard models. To address this issue, we proposed a novel over-defense data sampling framework and introduced the NotInject benchmark to systematically evaluate the extent of over-defense in target prompt guard models. Furthermore, we presented InjecAgent, an advanced prompt guard model trained using a novel robust training strategy, MOF, designed to effectively mitigate model bias toward specific words. Our experimental results on various benchmark demonstrate that InjecAgent significantly outperforms existing prompt guard models across multiple dimensions, with substantial improvements in performance and robustness. To the best of our knowledge, this is the first work to provide a fully open-source prompt guard model against injection, including the training dataset, strategies, code, and model. We believe this will further promote transparency and foster an open-source academic research environment for advancing future LLM safety exploration.
@@ B.1 Full Results
B.2 Visualization of NotInject dataset
@@
Appendix C Prompts
@@
C.1 Word-based Generation Prompt
@@ C.2 Refinement Prompt
C.3 Long-tail Format Augmentation Prompt
@@
Figure 8: The error ratio of three subsets after LLM generation.

### v3

@@ 5.1 Experimental Setups
Evaluation Datasets. We evaluate (1) benign accuracy, which utilizes benign data from both the PINT benchmark LakeraAI (2024b) and WildGuard benchmark Han et al. (2024); (2) malicious accuracy, which uses attack data from both the PINT benchmark LakeraAI (2024b) and BIPIA datasets Yi et al. (2023) and over-defense on our proposed NotInject dataset.
[…]
@@ 5.2 Main Results
We compare our method with other baselines in terms of performance and overhead. The results are shown in Tab. 1 (detailed results for each subset deferred to Appendix. B.1).
Performance Comparison. Our InjecGuard demonstrates superior performance compared to existing prompt guard models and even rivals commercial LLMs like GPT-4o. Specifically, InjecGuard achieves an average accuracy of 83.48%, the best-performing commercial prompt guard model, LakeraGuard, by 6.25%, and exceeding the top open-source model, ProtectAIv2, by 30.83%. Our model excels across all evaluation categories, attaining an over-defense accuracy of 87.32%, benign accuracy of 85.74%, and malicious accuracy of 77.39%. This balanced performance indicates that InjecGuard is highly effective at correctly identifying benign and malicious inputs while minimizing false positives and negatives. The fact that InjecGuard achieves results comparable to GPT-4o, despite being based on the open-source data and lightweight backbone, highlights the efficiency and effectiveness of our training approach. Notably, according to Tab. 1, the Efficiency (Performance/Inference time) of our model is 503 times higher than that of GPT-4o.
@@
Overhead Comparison. Our model, InjecGuard, excels in both computational efficiency and performance compared to existing baseline models. As shown in Tab. 1, InjecGuard  achieves an average accuracy of 83.48% while maintaining a GFLOPS count of 60.45 and an inference time of only 15.34 milliseconds. With an efficiency score of 5.44 (performance divided by inference time), InjecGuard delivers robust results swiftly and resource-effectively, making it ideal for environments prioritizing both accuracy and efficiency. In contrast, LLMs like GPT-4o achieve slightly higher accuracy (85.53%) but at a significant computational cost, with an inference time of 7907.18 milliseconds and an efficiency score of 0.01. Similarly, models such as Llama-2-chat and LlamaGuard3 require thousands of GFLOPS and longer inference times, resulting in much lower efficiency.
Training Design Overdef. Benign Mal. Avg.
Basic dataset training 75.22 78.53 70.17 74.64
+ Data-centric augment 64.31 81.36 75.95 73.87
Basic dataset training 75.22 78.53 70.17 74.64
+ MOF scratch retrain 89.38 84.73 71.57 81.89
Basic dataset training 75.22 78.53 70.17 74.64
+ Data-centric augment 64.31 81.36 75.95 73.87
+ MOF with finetuning 68.14 82.11 74.81 75.02
Basic dataset training 75.22 78.53 70.17 74.64
+ Data-centric augment 64.31 81.36 75.95 73.87
+ MOF scratch retrain 87.32 85.74 77.39 83.48
Table 2: Ablation study of each training design.
@@ 5.3 Ablation Studies
The ablation study presented in Tab. 2 offers key insights into the effects of our training components on InjecGuard performance. Starting with the Basic Dataset (Sec. 4.1), the model establishes a baseline average accuracy of 74.64%, with over-defense, benign, and malicious accuracies of 75.22%, 78.53%, and 70.17%, respectively. Introducing Data-centric Augmentation alone leads to a noticeable improvement in benign accuracy (from 78.53% to 81.36%) and malicious accuracy (from 70.17% to 75.95%), enhancing the model’s ability to correctly classify both benign and malicious inputs. However, this augmentation comes at a cost, as evidenced by a significant reduction in over-defense accuracy from 75.22% to 64.31%. This decline indicates that while data-centric augmentation enriches the training data and improves classification capabilities, it inadvertently exacerbates the over-defense issue, making the model more prone to incorrectly classifying benign inputs as malicious—a challenge commonly observed in existing models.
@@
Method Over-defense (%) Malicious (%)
InjecGuard (w/o MOF) 64.31 75.95
+ shortcut mitigation 86.73 65.53
+ MOF 87.32 77.39
Table 3: Comparison with shortcut mitigation method.
5.4 Comparison with Shortcut Mitigation
In Sec. 2, we discuss the susceptibility of injection attack detection tasks to spurious correlations. These correlations emerge from the interplay between semantic understanding and pattern recognition requirements, potentially challenging conventional shortcut mitigation methods. To further investigate the effectiveness of MOF compared to typical shortcut mitigation methods, we implement a representative shortcut mitigation approach Wang et al. (2021) to InjecGuard without MOF. The results presented in Tab. 3 show that while the shortcut mitigation method Wang et al. (2021) improves over-defense performance, it leads to a significant 10.42% decline in malicious performance, a well-known issue in current shortcut mitigation techniques Khani and Liang (2021); Mendelson and Belinkov (2021). In contrast, MOF not only achieves superior overall defense performance (87.32%) but also improves malicious performance to 77.39%, demonstrating the advantages of MOF over existing shortcut mitigation methods in this pattern-based task.
@@ 6 Conclusions
In this paper, we demonstrated the over-defense phenomenon in existing prompt guard models. To address this, we introduced the NotInject benchmark to evaluate the extent of over-defense. Furthermore, we presented InjecGuard, a prompt guard model that can significantly outperform existing models in both performance and robustness. To the best of our knowledge, this is the first work to provide a fully open-source prompt guard model against injection, including the training dataset, strategies, code, and model. We believe this will further promote transparency and foster an open-source academic research environment for advancing future LLM safety exploration.
@@ A.2 Malicious dataset
Category one-word two-word three-word
Common Queries 58 49 19
Techniques 16 30 41
Virtual Creation 14 4 24
Multilingual 25 30 29
Table 8: Topic category distribution of NotInject.
@@ B.1 Full results
Number over-defense benign malicious average
500 74.63 82.19 80.55 79.12
1,000 87.32 85.74 77.39 83.48
2,000 92.63 85.99 71.12 83.24
Table 9: Ablation study of MOF sampling scale.
B.2 Ablation study of MOF sampling scale
In Sec. 4.2, MOF generates 1,000 benign samples containing biased tokens, which are incorporated into the training process to reduce the model’s over-defense degree. To explore the optimal scale for generating benign samples, we conduct experiments using MOF to generate samples at varying scales and evaluate the performance of InjecGuard trained on these datasets. The results shown in Tab. 9 reveal a clear trade-off between over-defense and malicious accuracy as the number of generated samples increases. Notably, training with 1,000 generated samples achieves the best balance, resulting in the highest average accuracy of 83.48%. We thereby select 1,000 as the sampling scale.
B.3 The effect of MOF on general over-defense scenarios
Although MOF primarily focuses on non-trigger word scenarios, we further explore the effectiveness of MOF in general over-defense scenarios. For the experimental details, since there is no existing benchmark for non-trigger-word-based over-defense, we adopt the hard-negative dataset from the PINT LakeraAI (2024b) benchmark as a substitute, which can somewhat reflect the model’s performance in non-trigger-word-based over-defense scenarios.
The results in Tab. 10 clearly show that incorporating MOF improves performance on the hard-negative dataset, demonstrating its effectiveness in mitigating over-defense in non-trigger-word-based scenarios. The accuracy of 91.15% further highlights the robustness of InjecGuard across general over-defense scenarios.
Method hard negative accuracy
InjecGuard (w/o MOF) 87.86
InjecGuard (w/ MOF) 91.15
Table 10: The effect of MOF on general over-defense scenarios.
Appendix C Qualitative Analysis
C.1 Visualization of NotInject dataset
@@
C.2 The robustness of MOF when incorrectly identified words
In this section, we explore the robustness of the MOF when trigger words are incorrectly identified and assess whether such misidentifications could harm model performance. In fact, incorrectly identified words do not have a detrimental effect on the model. This is because these misidentified words are used to construct benign sentences, which do not introduce noisy labels or degrade the model’s learning process. For example, if an unbiased word like “book” is mistakenly identified as biased, and we construct a sentence such as “Can you recommend some history books for me?” adding this sentence to the training dataset would merely serve as a benign augmentation. While this might slightly reduce the degree of improvement in over-defense performance, it would not negatively impact the model’s basic performance.
Figure 6: The human curation agreement in trigger word refinement process.
@@ C.2 The robustness of MOF when incorrectly identified words
Figure 8: Over-denfese issue in PromptGuard Meta (2024) 1.
Figure 10: The error ratio of three subsets after LLM generation.
Algorithm 1 Trigger Words Identification
1:Benign dataset $D_{b}$, Malicious dataset $D_{m}$, Integer $k$
2:Compute word frequencies in $D_{b}$ to get frequency list $F_{b}$
3:Compute word frequencies in $D_{m}$ to get frequency list $F_{m}$
4:Sort $F_{b}$ in descending order to get rank list $R_{b}$
5:Sort $F_{m}$ in descending order to get rank list $R_{m}$
6:for all words $w$ in $R_{b}\cup R_{m}$ do
7:     Compute rank difference $(\Delta r(w)=R_{b}(w)-R_{m}(w))$
8:     Add $(w,\Delta r(w))$ to list $L$
9:end for
10:Sort list $L$ in descending order based on $\Delta f(w)$
11:return Top $k$ words from list $L$
Appendix D Prompts
@@
D.1 Word-based Generation Prompt
@@ D.2 Refinement Prompt
D.3 Long-tail Format Augmentation Prompt
@@
11footnotetext: https://huggingface.co/meta-llama/Prompt-Guard-86M